OpenStudy (anonymous):

Please help! I don't have any idea what to do! Will fan and medal! Suppose f(x) is a polynomial of degree 4 or greater such that f(1)=2, f(2)=3, and f(3)=5. Find the remainder when f(x) is divided by (x-1)(x-2)(x-3).

jimthompson5910 (jim_thompson5910):

Does this page help at all? http://mathforum.org/library/drmath/view/60310.html look at problem (2)

OpenStudy (anonymous):

It helps a little but I am still unable to solve the question. Could you explain further?

jimthompson5910 (jim_thompson5910):

what is the degree of (x-1)(x-2)(x-3) ? Any idea? hint: expand out the polynomial

OpenStudy (anonymous):

degree 3

jimthompson5910 (jim_thompson5910):

so the degree of the remainder is going to be less than 3 http://www.sosmath.com/algebra/factor/fac01/fac01.html since the degree of the remainder is always less than the degree of the divisor

jimthompson5910 (jim_thompson5910):

so the remainder after dividing by (x-1)(x-2)(x-3) will either be a single number OR something in the form Ax+B OR something in the form Ax^2 + Bx + C hopefully you agree so far?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

Let's assume it's in the form Ax^2 + Bx + C I'm going to use the same basic approach as the professor is showing on the first link I posted If that's the case, then f(x)/[(x-1)(x-2)(x-3)] = Q(x) + R(x)/[(x-1)(x-2)(x-3)] which turns into f(x) = (x-1)(x-2)(x-3)*Q(x) + R(x) f(x) = (x-1)(x-2)(x-3)*Q(x) + Ax^2 + Bx + C

jimthompson5910 (jim_thompson5910):

Now we plug in x = 1 f(x) = (x-1)(x-2)(x-3)*Q(x) + Ax^2 + Bx + C f(1) = (1-1)(1-2)(1-3)*Q(1) + A*(1)^2 + B*(1) + C f(1) = A+B+C 2 = A+B+C A+B+C = 2

jimthompson5910 (jim_thompson5910):

next plug in x = 2 f(x) = (x-1)(x-2)(x-3)*Q(x) + Ax^2 + Bx + C f(2) = (2-1)(2-2)(2-3)*Q(2) + A(2)^2 + B(2) + C 3 = 4A + 2B + C 4A + 2B + C = 3

jimthompson5910 (jim_thompson5910):

and finally x = 3 f(x) = (x-1)(x-2)(x-3)*Q(x) + Ax^2 + Bx + C f(3) = (3-1)(3-2)(3-3)*Q(3) + A(3)^2 + B(3) + C 5 = 9A + 3B + C 9A + 3B + C = 5

OpenStudy (anonymous):

so I should just solve the system of equations?

jimthompson5910 (jim_thompson5910):

so we have this system of equations A+B+C = 2 4A + 2B + C = 3 9A + 3B + C = 5 are you able to solve for A,B,C ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So I can add the first 2 equations to get 5A+3B+2C=5 and then subtract that from the bottom equation to get: 4A-C=0 or 4A=C.

jimthompson5910 (jim_thompson5910):

so because C = 4A, we can replace every copy of C with 4A to go from this system A+B+C = 2 4A + 2B + C = 3 9A + 3B + C = 5 to this system A+B+4A = 2 4A + 2B + 4A = 3 9A + 3B + 4A = 5 pick any two equations you want and solve for A,B

OpenStudy (anonymous):

So the second equation becomes 2C+2B=3 or C+B=1.5 So A=.5, B=-.5, and C=2.

jimthompson5910 (jim_thompson5910):

those are the correct values for A,B,C

jimthompson5910 (jim_thompson5910):

so IF the remainder is of degree 2, then the remainder is the function R(x) = 0.5x^2 - 0.5x + 2

OpenStudy (anonymous):

Okay. Thanks for all of your help. I'll just check to make sure that is the correct answer.

jimthompson5910 (jim_thompson5910):

no problem